Already have an account? Get multiple benefits of using own account!
Login in your account..!
Remember me
Don't have an account? Create your account in less than a minutes,
Forgot password? how can I recover my password now!
Enter right registered email to receive password!
What is Conditional Probability?
Two events A and B are said to be dependent when B can occur only when A is known to have occurred (or vice versa). The probability attached to such an event is called the conditional probability and is denoted byP(A/B) or, in other words probability of A given that B has occurred.If two events A and B are dependent, then the conditional probability of B given A is:P(B/A) = P(AB)/P(A)Proof: suppose a1 is the number of cases for the simultaneous happening of A and B out of a1 + a2 cases in which A can happen with or without happening of B.∴ P(B/A) = a1/(a1 + a2) = (a1/n)/(a1 + a2)/n = P(AB)/P(A)Similarly it can be shown thatP(A/B) = P[(AB)/P(B)]The general rule of multiplication in its modified form in terms of conditional probability becomes:P(A and B) = P(B) × P(A/B)Or, P(A and B) = P(A) × P(B/A)For three events A, B and C we haveP(ABC) = P(A) × P(B/A) × P(C/AB)i.e. the probability of occurrence of A, B and C is equal to the probability of A, times of the probability of B given that A has occurred, times the probability of C given that both A and B have occurred.Illustration: a bag contains 5 white and 3 black balls. Two balls are drawn at random one after the other without replacement. Find the probability that both balls drawn are black.Solution: probability of drawing a black ball in the first attempt isP(A) = 3/(5 + 3) = 3/8Probability of drawing the second black ball given that the first ball drawn is blackP(B/A) = 2/(5 + 2) = 2/7∴ The probability that both balls drawn are black is given byP(AB) = P(A) × P(B/A) = 3/8 × 2/7 = 3/28
Math Assignment Help
Question: (a) Explain the purpose of internal marketing. (b) Identify the three levels organizational culture and discuss the importance organizational culture for successf
What is the working of final or last barrier to exchange in channels distribution? All problem are signifies the discrepancy of quantity and assortment within the exchange proc
WV8 is a large not for profit (NFP) public sector organisation responsible for organising the operation of publicly financed schools within its region. WV8 receives its income sole
value pricing products
considering the concept of product life cycle, where would you put video games in their life cycle
briefly expain trend describe the impact of the trend on the idustry
The balance of the communications mix Role of advertising and sales promotions in business-to-business communications is mainly to support personal selling effort. This cont
What is market segmentation? Market Segmentation: Market segmentation is the process of dividing a total market into group of consumers who have relatively similar produc
A curve which touches each member of a family of curves and which is touched at each point by some member of the family is called the envelope of the family of curves. Example: Co
Boolean algebra differs from normal algebra in three different ways: The values are of a logical and not of a numerical character. The operations applicable to those values. The pr
Get guaranteed satisfaction & time on delivery in every assignment order you paid with us! We ensure premium quality solution document along with free turntin report!
whatsapp: +91-977-207-8620
Phone: +91-977-207-8620
Email: [email protected]
All rights reserved! Copyrights ©2019-2020 ExpertsMind IT Educational Pvt Ltd